OpenStudy (anonymous):

Solve the equation. Check both solutions and only write the real solution. the square root of 6m minus 5 = m A. {1} B. {5} C. {1, 5} D. no real solution

OpenStudy (anonymous):

\[\sqrt{6m-5=m}\]

OpenStudy (turingtest):

\[\sqrt{6m-5}=m\] you mean

OpenStudy (anonymous):

Since it gives us the answers, let's guess and check! \[\sqrt{6(1)-5}=\sqrt{1}=1\]\[\sqrt{6(5)-5}=\sqrt{25}=5\] So it is both [1,5]

OpenStudy (anonymous):

\[\sqrt{6m-5=m}\]......\[\sqrt{5m=5}\]....\[\sqrt{1}=1\]\[\sqrt{25}=5\]....both 1 and 5